OpenStudy (anonymous):

find the composition fog of the functions. f(x)=x+5 over 2x+1 g(x)=2x+1 over x+5 please help got stuck in this one

OpenStudy (anonymous):

fOg(x)= f(g(x)) f(x)= \[\frac{ x+5 }{2x+1 }\] g(x) = \[\frac{ 2x+1}{x+5 }\] 

OpenStudy (anonymous):

In the last step it is 4x+2+x+5 in the denominator. So it is 7x+26 over 5x+7.